OpenStudy (anonymous):

I do not know what the freak my professor assigned....this "make the subject" is not working out. Make "a" the subject of c = sqrt of a^2 - b^2 (yes, its a squared minus b squared in the lil square root thingy)

OpenStudy (anonymous):

when they ask that they mean to define the equation in terms of "whatever variable"

OpenStudy (anonymous):

so now solve the equation in terms of "a"

myininaya (myininaya):

\[c=\sqrt{a^2-b^2}\] \[c^2=a^2-b^2\] \[c^2+b^2=a^2\] \[\pm \sqrt{c^2+b^2}=a\]

OpenStudy (anonymous):

Yo u r a beast myiniaya ! Thanks.

myininaya (myininaya):

lol

OpenStudy (anonymous):

\[\color{green}{\text{jealous}}\]